# Age-Optimal Sampling and Routing under Intermittent Links and Energy Constraints

### v1

@@ Abstract
Links in practical systems, such as satellite–terrestrial integrated networks, exhibit distinct delay distributions, intermittent availability, and heterogeneous energy costs. These characteristics pose significant challenges to maintaining timely and energy-efficient status updates. While link availability restricts feasible transmission routes, routing decisions determine the actual delay and energy expenditure. This paper tackles these challenges by jointly optimizing sampling and routing decisions to minimize monotonic, non-linear Age of Information (AoI). The proposed formulation incorporates key system features, including multiple routes with correlated random delays, stochastic link availability, and route-dependent energy consumption. We model the problem as an infinite-horizon Constrained Semi-Markov Decision Process (CSMDP) with a hybrid state–action space and develop an efficient nested algorithm, termed Bisec-ReaVI, to solve this problem. We reveal a well-defined jointly optimal policy structure: (i) The optimal routing policy is a monotonic handover policy that adapts to the availability of routes and their mean delays; and (ii) The optimal sampling policy is a piecewise linear waiting policy, with at most $\binom{N}{2}+N$ breakpoints given $N$ routes. Numerical experiments in a satellite–terrestrial integrated routing scenario demonstrate that the proposed scheme efficiently balances energy usage and information freshness, and reveal a counter-intuitive insight: even routes with higher average delay, higher delay variance or lower availability can still play a critical role in minimizing monotonic functions of AoI.
@@ I-A Background
This growing emphasis on information freshness has led to the development of the Age of Information (AoI) metric that quantifies it [kaul2012real]. Distinct from traditional latency, AoI provides a receiver-centric, flow-level measure of information timeliness. Formally, at any time $t$, the AoI is defined as $\Delta(t)\triangleq t-U(t)$ , where $U(t)$ denotes the generation time of the latest received sample. Maintaining a low AoI requires both sufficiently frequent updates and low-latency delivery, thus coupling throughput and delay in a novel performance metric. Hence, AoI combines the conventional performance metrics of latency and throughput in a novel way. Over the past few years, AoI minimization has been studied under various constraints and network models, including energy-limited update policies [bacinoglu2015age, yates2015lazy, bacinoglu2019optimal], multi-hop and multi-source networks [talak2017minimizing, beytur2018minimizing, kadota2019scheduling, bedewy2021optimal, ceran2021reinforcement], broadcast settings [kadota2018scheduling], and unreliable or error-prone communication channels [7541765, 9931172, 11184587] (see [DBLP:journals/jsac/YatesSBKMU21a] for a comprehensive review). Furthermore, in scenarios where the significance of data extends beyond temporal freshness, more sophisticated evaluation frameworks have been developed [chen2022uncertainty, chen2024optimal, sun2019sampling, 10806969, 10807024, shisher2024timely, sun2019sampling2]. Some of these frameworks utilize AoI as an intermediate metric to capture task-specific relevance through the freshness of data samples [shisher2024timely].
@@ I-B Motivation
In this paper, we focus on the optimization of AoI. Our goal is to extend the formulation of Age-optimal sampling first proposed in [sun2017update] to a case where there are multiple routing options between the source and the destination. The new formulation proposed in this paper was inspired and motivated by the growing interest in Satellite IoT and integrated TN-NTN in 5G and 6G, where data transmission decisions are sometimes faced with a choice between routing through non-terrestrial links versus terrestrial connections:
@@ I-C Related Works and Contributions
  • System Model: We formulate a joint sampling and routing problem in which a transmitter optimizes both the sampling interval and the transmission route to minimize the long-term average AoI at the destination, subject to an average routing energy constraint and route availability constraint. Our formulation is a direct extension of the problem in [sun2017update], and in contrast to this and other prior work where the sampling problem is attacked under given delay statistics [sun2019sampling, 10004990, 10806969, 10807024], our system model takes a proactive approach by actively selecting and switching routes to control the delay experienced by status updates. Meanwhile, unlike existing multi-channel scheduling problems that either focus on homogeneous channels with uniform one-slot delays [sun2023age] or heterogeneous channels where each channel experiences an on-off constant discrete delay [pan2021minimizing, 9882370], our work considers distinct continuous delay distributions across different routes and jointly optimizes sampling and route selection strategies. This works generalizes our prior work in [Mobihoc25] by incorporating the following practical aspects: (i) Correlated route delays, motivated by shared infrastructures or congestion effects that couple delay statistics across different routes; (ii) Stochastic route availability, modeled as independent random on/off links with fixed availability probabilities, capturing real-world uncertainties such as satellite visibility, atmospheric attenuation, and space-weather-induced disruptions and (iii) Route-dependent energy consumption under an average energy constraint, which is essential for energy-limited platforms such as satellites or IoT devices.
@@
  • Solution Methodology: We show the problem can be formulated by a Constrained Semi-Markov Decision Process (CSMDP) with uncountable hybrid state and action spaces. The state captures discrete link availability and continuous delay, while the action includes both routing choices and sampling intervals. Such CSMDPs are known to be challenging to solve due to the size of the state and action spaces, and previous research has addressed this complexity by: ($i$) discretizing the uncountable state space and the action space [8712546, zhou2024age], which introduces quantization error or; ($ii$) focusing on a special case of SMDP where the state transitions are independent of actions [^1] This allows the Markov decision process to be reduced to a renewal reward process, thus simplifying the analysis. [sun2019sampling, sun2017update, 10004990], which, however, does not hold in our scenario. In this paper, we develop a new nested algorithm namely Bisection Relative Expected Action Value Iteration (Bisec-ReaVI) that directly solves this class of SMDPs without discretizing the space. To the best of our knowledge, this is the first algorithm that efficiently solves hybrid-state CSMDPs while preserving structural optimality and avoiding discretization error.
@@
  • Structural Results: We prove that each of the jointly optimal sampling and routing policies exhibit a graceful threshold structure: (ii) The routing policy is a thresholds-based handover policy, where a specific route is selected when the current AoI at the receiver falls within certain range, precisely determined by multiple thresholds; ($ii$) A new sample is taken and transmitted when the AoI at the receiver reaches a threshold that is a function of the selected transmission route. These structural properties deem the policies suitable for practical implementation. We designed an efficient algorithm to determine these thresholds.
@@
  • Counter-Intuitive Insights: We test our algorithms on the model of an integrated satellite-terrestrial communication network scenario. Our simulation results reveal an intriguing insight: routes with higher mean delay, greater variance, or lower availability can still contribute to minimizing AoI. This finding challenges conventional wisdom that may prioritize routes with minimal mean delay or delay variance characteristics. It demonstrates that the strategic utilization of diverse routing options in complex network environments can lead to superior information freshness.
@@ II-A Persistent and Intermittent Links
In this work, we consider a heterogeneous network consisting of multiple communication routes, denoted collectively as the set $\mathcal{N}$. These routes are categorized into two disjoint subsets based on their physical characteristics and long-term availability:
  • Persistent Routes ($\mathcal{R}_{\infty}$): The routes in set $\mathcal{R}_{\infty}$ are continuously accessible over time and typically correspond to terrestrial links such as fiber-optic or cellular infrastructure. Due to their stable physical environment and minimal susceptibility to external disruptions, these links exhibit deterministic availability, and are modeled with an availability probability of $p_{k}=1$. The set of persistent routs are given as:
    $\mathcal{R}_{\infty}=\{k\in\mathcal{R}:p_{k}=1\}\subset\mathcal{N}.$ (1)
    To guarantee continuous data transmission, we assume that $\mathcal{R}_{\infty}\neq\emptyset$, which corresponds to the presence of at least one persistently available terrestrial route (e.g., cellular or fiber-optic), ensuring baseline connectivity even when all intermittent routes are unavailable.
  • Intermittent Routes ($\mathcal{R}_{<\infty}$): This subset includes routes whose availability varies over time due to stochastic physical factors. Typical examples include satellite links or other opportunistic channels affected by satellite orbital motion, line-of-sight (LOS) constraints, or environmental interference (e.g., weather conditions). Each intermittent route $k\in\mathcal{R}_{<\infty}$ is characterized by a stationary availability probability $p_{k}\in(0,1)$, which denotes the long-term fraction of time the route is usable. The set of persistent routs are given as:
    ${R}_{<\infty}=\{k\in\mathcal{R}:0<p_{k}<1\}\subset\mathcal{N}.$ (2)
@@ II-C Heterogeneous Energy Costs
The system is subject to a long-term average energy constraint denoted by $E_{\max}$, which limits the energy consumption over time. This constraint is particularly relevant in energy-constrained systems such as remote sensing applications or satellite-terrestrial networks, where power sources (e.g., battery-powered ground terminals or solar-powered satellite relays) are limited. Each update cycle is associated with two distinct types of energy expenditures:
@@
We assume a non-preemptive system, where a new transmission can begin only after the previous one has been completed [sun2017update]. Upon receiving each data sample, the monitor sends an ideal acknowledgment (ACK) to the transmitter, indicating that the system is ready to initiate the next transmission.
We adopt the generate-at-will model [sun2017update, yates2015lazy], in which the sampler can become active at any time, provided that a new transmission is allowed. We next introduce some notation. After receiving the ACK corresponding to the $i$-th transmission, the $(i+1)$-th data sample is generated and submitted to route $R_{i}$, selected from the pool of available routes $\mathcal{R}_{i}$, at time instant $S_{i+1}$. It is subsequently delivered to the monitor at time instant $D_{i+1}=S_{i+1}+Y_{{i+1},R_{i}}$}. The overall energy cost for transmitting the $(i+1)$-th update is:
$E_{i+1}=C_{s}+G_{R_{i}}\dot{Y}_{i+1,R_{i}}.$ (5)
II-D Age of Information
The Age of Information (AoI) is the metric of our interest to measure the freshness of information. This metric is defined as the time elapsed since the generation of the most recently received data sample [kaul2012real]. Specifically, the AoI $\Delta(t)$ at time $t$ is defined by [^2] The framework in this work can be readily extended to general AoI penalty functions $g(\Delta(t))$ as alternative metrics of interest, where $g(\cdot):\mathbb{R}^{+}\rightarrow\mathbb{R}^{+}$ represents any monotonically non-decreasing function. However, due to page limits, we focus on the linear case in this paper.
@@ III Problem Formulation
We aim at minimizing the long-term average AoI by designing a joint sampling and routing policy $\pi\triangleq(R_{0},Z_{0},R_{1},Z_{1}\dots)$. This policy consists of two distinct sequences: (i) a sequence of routing decisions $(R_{0},R_{1},R_{2},\dots)$, where $R_{i}$ specifies the route selected for transmitting the $(i+1)$-th packet, and (ii) a sequence of finite waiting times $(Z_{0},Z_{1},Z_{2},\dots)$, where $Z_{i}<\infty$ determines the $(i+1)$-th sampling (or submission) time as $S_{i+1}=D_{i}+Z_{i}$. Let $\Pi$ denote the set of all causal and stationary policies $\pi$. WE consider the following problem:
Problem 1 (Average Age Minimization with Energy Constraint).
$\lambda^{\star}=$ $\min_{\pi\in\Pi}\;\limsup_{T\to\infty}\;\frac{1}{T}\mathbb{E}\!\left[\int_{0}^{T}\Delta(t)\,dt\right]$ (7)
@@
where $N_{s}(T)$ is the number of sampling actions up to time $T$, and $\lambda^{\star}$ is the optimal long-term average AoI. This problem aims at minimizing the long-term average age under a long-term average energy constraint.
Figure 2: Sample evolution of the AoI process $\Delta(t)$.
@@
$\limsup_{T\to\infty}\mathbb{E}\!\left[\frac{1}{T}\int_{0}^{T}\!\Delta(t)\,dt\right]=$ (8)
$\limsup_{n\to\infty}\frac{\mathbb{E}\!\left[\frac{1}{n}\sum_{i=0}^{n-1}q(Y_{i,R_{i-1}},Y_{i+1,R_{i}},Z_{i})\right]}{\mathbb{E}\!\left[\frac{1}{n}\sum_{i=0}^{n-1}(Z_{i}+Y_{i+1,R_{i}})\right]},$
where $q(Y_{i,R_{i-1}},Y_{i+1,R_{i}},Z_{i})$ represents the accumulated AoI (the area under $\Delta(t)$) during the $i$-th cycle.
In addition, the energy process $E(t)$ is also a piecewise constant function. Similarlly, the long-term average energy consumed up to the $n$-th update can be written as:
@@
$=\limsup_{n\to\infty}\frac{\mathbb{E}\!\left[\frac{1}{n}\sum_{i=0}^{n-1}E_{i+1}\right]}{\mathbb{E}\!\left[\frac{1}{n}\sum_{i=0}^{n-1}(Z_{i}+Y_{i+1,R_{i}})\right]},$
@@
Problem 2 (Reformulated Average Age Minimization).
$\lambda^{\star}\triangleq\min_{\pi}\;$ $\limsup_{n\to\infty}\;\frac{\mathbb{E}\!\left[\frac{1}{n}\sum_{i=0}^{n-1}q(Y_{i,R_{i-1}},Y_{i+1,R_{i}},Z_{i})\right]}{\mathbb{E}\!\left[\frac{1}{n}\sum_{i=0}^{n-1}(Z_{i}+Y_{i+1,R_{i}})\right]}$ (10a)
s.t. $\limsup_{n\to\infty}\;\frac{\mathbb{E}\!\left[\frac{1}{n}\sum_{i=0}^{n-1}E_{i+1}\right]}{\mathbb{E}\!\left[\frac{1}{n}\sum_{i=0}^{n-1}(Z_{i}+Y_{i+1,R_{i}})\right]}\;\leq\;E_{\max}$ (10b)
Problem 2 is an infinite-horizon Constrained Semi-Markov Decision Process (CSMDP) with a hybrid state space and a state-dependent hybrid action space, posing significant challenges due to its fractional objective and the presence of a long-term energy constraint. To tackle this, we adopt a divide-then-concur strategy, consisting of three key steps: $i$) Fractional constraint reformulation: We first transform the original fractional constraint into an equivalent linear form to simplify the constraint structure; $ii$) Dinkelbach transformation: Next, we apply the Dinkelbach approach [dinkelbach1967nonlinear] to convert the fractional objective into a sequence of parameterized CMDP problems; iiiiii) Lagrangian techniques: Finally, we employ the Lagrangian method to handle the constraint in the Constrained Markov Decision Process (CMDP) formulation, effectively reducing the problem to a standard MDP.

### v2

@@ Abstract
Links in practical systems, such as satellite–terrestrial integrated networks, exhibit distinct delay distributions, intermittent availability, and heterogeneous energy costs. These characteristics pose significant challenges to maintaining timely and energy-efficient status updates. While link availability restricts feasible transmission routes, routing decisions determine the actual delay and energy expenditure. This paper tackles these challenges by jointly optimizing sampling and routing decisions to minimize monotonic, non-linear Age of Information (AoI). The proposed formulation incorporates key system features, including multiple routes with correlated random delays, stochastic link availability, and route-dependent energy consumption. We model the problem as an infinite-horizon Constrained Semi-Markov Decision Process (CSMDP) with a hybrid state–action space and develop an efficient nested algorithm, termed Bisec-ReaVI, to solve this problem. We analyze the structural properties of the solution and reveal a well-defined jointly optimal policy structure: (i) For general monotonic penalty functions, the optimal sampling policy is a piecewise linear waiting policy with at most $N$ breakpoints given $N$ routes; and (ii) under a derived Expected Penalty Ordering condition, the optimal routing policy is a monotonic threshold-based handover policy characterized by at most $\binom{N}{2}$ thresholds. Numerical experiments in a satellite–terrestrial integrated routing scenario demonstrate that the proposed scheme efficiently balances energy usage and information freshness, and reveal a counter-intuitive insight: even routes with higher average delay, higher delay variance or lower availability can still play a critical role in minimizing monotonic functions of AoI.
@@ I-A Background
This growing emphasis on information freshness has led to the development of the Age of Information (AoI) metric that quantifies it [16]. Distinct from traditional latency, AoI provides a receiver-centric, flow-level measure of information timeliness. Formally, at any time $t$, the AoI is defined as $\Delta(t)\triangleq t-U(t)$, where $U(t)$ denotes the generation time of the latest received sample. Maintaining a low AoI requires both sufficiently frequent updates and low-latency delivery, thus coupling throughput and delay in a novel performance metric. Hence, AoI combines the conventional performance metrics of latency and throughput in a novel way. Over the past few years, AoI minimization has been studied under various constraints and network models, including energy-limited update policies [3, 35, 4], multi-hop and multi-source networks [32, 6, 21, 14, 5, 18, 7], broadcast settings [15], and unreliable or error-prone communication channels [9, 19, 36] (see [34] for a comprehensive review). Furthermore, in scenarios where the significance of data extends beyond temporal freshness, more sophisticated evaluation frameworks have been developed [8, 11, 29, 20, 10, 26, 27]. Some of these frameworks utilize AoI as an intermediate metric to capture task-specific relevance through the freshness of data samples [26].
@@ I-B Motivation
In this paper, we focus on the optimization of monotonic functions of AoI. Our goal is to extend the formulation of Age-optimal sampling first proposed in [31] to a case where there are multiple routing options between the source and the destination. The new formulation proposed in this paper was inspired and motivated by the growing interest in Satellite IoT and integrated TN-NTN in 5G and 6G, where data transmission decisions are faced with choosing between routing through non-terrestrial links versus terrestrial connections:
@@ I-C Related Works and Contributions
  • System Model: We formulate a joint sampling and routing problem in which a transmitter optimizes both the sampling interval and the transmission route to minimize the long-term average of a monotonic function of the AoI at the destination, subject to an average routing energy constraint and route availability constraint. Our formulation is a direct extension of the problem in [31], and in contrast to this and other prior work where the sampling problem is attacked under given delay statistics [29, 33, 20, 10], our system model takes a proactive approach by actively selecting and switching routes to control the delay experienced by status updates. Meanwhile, unlike existing multi-channel scheduling problems that either focus on homogeneous channels with uniform one-slot delays [28] or heterogeneous channels where each channel experiences an on-off constant discrete delay [22, 23], our work considers distinct continuous delay distributions across different routes and jointly optimizes sampling and route selection strategies. This work generalizes our prior work in [2] by incorporating the following practical aspects: (i) Correlated route delays, motivated by shared infrastructures or congestion effects that couple delay statistics across different routes; (ii) Stochastic route availability, modeled as independent random on/off links with fixed availability probabilities, capturing real-world uncertainties such as satellite visibility, atmospheric attenuation, and space-weather-induced disruptions; (iii) Route-dependent energy consumption under an average energy constraint, which is essential for energy-limited platforms such as satellites or IoT devices; and (iv) Non-linear age penalty functions, extending the optimization framework beyond standard linear AoI to support general monotonic functions of age, allowing for more flexible service requirement modeling.
@@
  • Solution Methodology: We show the problem can be formulated by a Constrained Semi-Markov Decision Process (CSMDP) with uncountable hybrid state and action spaces. The state captures discrete link availability and continuous delay, while the action includes both routing choices and sampling intervals. Such CSMDPs are known to be challenging to solve due to the size of the state and action spaces, and previous research has addressed this complexity by: ($i$) discretizing the uncountable state space and the action space [17, 37], which introduces quantization error or; ($ii$) focusing on a special case of SMDP where the state transitions are independent of actions [^1] This allows the Markov decision process to be reduced to a renewal reward process, thus simplifying the analysis. [29, 31, 33], which, however, does not hold in our scenario. In this paper, we develop a new nested algorithm namely Bisection Relative Expected Action Value Iteration (Bisec-ReaVI) that directly solves this class of SMDPs without discretizing the space. To the best of our knowledge, this is the first algorithm that efficiently solves hybrid-state CSMDPs while preserving structural optimality and avoiding discretization error.
@@
  • Structural Results: We prove that each of the jointly optimal sampling and routing policies exhibit a graceful threshold structure: ($i$) The routing policy is a thresholds-based handover policy, where a specific route is selected when the current AoI at the receiver falls within certain range (a route may be optimal for more than one such regions), precisely determined by multiple thresholds; ($ii$) A new sample is taken and transmitted when the AoI at the receiver reaches a threshold that is a function of the selected transmission route. These structural properties deem the policies suitable for practical implementation.
@@
  • Counter-Intuitive Insights: We test our algorithms on the model of an integrated satellite-terrestrial communication network scenario. Our simulation results reveal an intriguing insight: routes with higher mean delay, greater variance, or lower availability can still contribute to minimizing AoI. This finding challenges conventional wisdom that may prioritize routes with minimal mean delay or delay variance characteristics. It demonstrates that the strategic utilization of diverse routing options can lead to superior information freshness.
@@ II-A Persistent and Intermittent Links
In this work, we consider a heterogeneous network consisting of multiple communication routes, denoted collectively as the set $\mathcal{N}$. These communication routes are categorized into two disjoint subsets based on their long-term availability:
  • Persistent Routes ($\mathcal{R}_{\infty}$): The routes in set $\mathcal{R}_{\infty}$ are continuously accessible over time and typically correspond to terrestrial links such as fiber-optic or cellular infrastructure. Due to their stable physical environment and minimal susceptibility to external disruptions, these links exhibit deterministic availability, and are modeled with an availability probability of $p_{k}=1$. The set of persistent routs are given as: $\mathcal{R}_{\infty}=\{k\in\mathcal{R}:p_{k}=1\}\subset\mathcal{N}.$ To guarantee continuous data transmission, we assume that $\mathcal{R}_{\infty}\neq\emptyset$, which corresponds to the presence of at least one persistently available terrestrial route (e.g., cellular or fiber-optic), ensuring baseline connectivity even when all intermittent routes are unavailable.
  • Intermittent Routes ($\mathcal{R}_{<\infty}$): This subset includes routes whose availability varies over time due to stochastic physical factors. Typical examples include satellite links or other opportunistic channels affected by satellite orbital motion, line-of-sight (LOS) constraints, or environmental interference (e.g., weather conditions). Each intermittent route $k\in\mathcal{R}_{<\infty}$ is characterized by a stationary availability probability $p_{k}\in(0,1)$, which denotes the long-term fraction of time the route is usable. The set of persistent routs are given as: ${R}_{<\infty}=\{k\in\mathcal{R}:0<p_{k}<1\}\subset\mathcal{N}.$
@@ II-B Correlated Random Delays
Remark 1 (Sufficiency of Marginals).
Although the delay vectors $\mathbf{Y}_{i}$ may exhibit correlation across routes within the same epoch (e.g., due to shared weather events affecting multiple satellite links), the optimal control policy depends exclusively on the marginal distributions $Q_{k}$. Since the system state resets upon delivery (due to the i.i.d. nature of epochs) and the decision maker observes only the realization of the selected route, the joint correlation structure does not provide exploitable information for future epochs. Thus, the optimization problem reduces to one defined purely by the set of marginals $\{Q_{k}\}_{k=1}^{N}$, and the potential correlation between unselected routes does not impact the achievable optimal age.
In the subsequent analysis, where the specific route index is determined by the policy, we will omit the subscript $k$ and denote the delay incurred during the $i$-th epoch simply as $Y_{i}$.
@@ II-C Heterogeneous Energy Costs
The system is subject to a long-term average energy constraint denoted by $E_{\max}$, which limits the energy consumption over time. This constraint is particularly relevant in energy-constrained systems such as remote sensing applications or satellite-terrestrial networks, where power sources (e.g., battery-powered ground terminals or solar-powered satellite relays) are limited. There are two types of energy expenditures:
@@
We assume a non-preemptive system, where a new transmission can begin only after the previous one has been completed [31]. Upon receiving each data sample, the monitor sends an ideal acknowledgment (ACK) to the transmitter, indicating that the system is ready to initiate the next transmission.
We adopt the generate-at-will model [31, 35], in which the sampler can become active at any time, provided that a new transmission is allowed. We next introduce some notation. After receiving the ACK corresponding to the $i$-th transmission, the $(i+1)$-th data sample is generated and submitted to route $R_{i}$, selected from the pool of available routes $\mathcal{R}_{i}$, at time instant $S_{i+1}$. It is subsequently delivered to the monitor at time instant $D_{i+1}=S_{i+1}+Y_{i+1}$}. The overall energy cost for transmitting the $(i+1)$-th update is:
$E_{i+1}=C_{s}+G_{R_{i}}Y_{i+1}.$ (3)
II-D Age of Information and No-linear Age
The Age of Information (AoI) is the metric of our interest to measure the freshness of information. This metric is defined as the time elapsed since the generation of the most recently received data sample [16]. Specifically, the AoI $\Delta(t)$ at time $t$ is defined by
@@ II-D Age of Information and No-linear Age
While the standard linear AoI metric captures the timeliness of updates, many practical systems exhibit sensitivity to information staleness that is non-linear. To address this, we consider a generalized age penalty function $f(\Delta(t))$, where $f(\cdot):[0,\infty)\to[0,\infty)$ is a continuous, monotonic non-decreasing function. This formulation allows us to model diverse application requirements: the standard linear case corresponds to $f(\Delta)=\Delta$, while scenarios demanding stricter freshness constraints can be modeled using exponential penalties and applications with saturation effects can be captured by logarithmic functions.
Furthermore, to ensure the problem is well-posed and the expected costs are finite, we adopt the following assumption on the penalty function. For any route $k\in\mathcal{N}$, the delay distribution $Q_{k}$ satisfies the finite expected penalty condition: $\mathbb{E}_{Y\sim Q_{k}}[f(Y)]<\infty.$
@@ III Problem Formulation
Figure 2: Sample evolution of the AoI process $\Delta(t)$.
Figure 3: Sample path of the non-linear $f(\Delta(t))=e^{0.35\Delta(t)}-1$.
We aim at minimizing the long-term average penalty of AoI by designing a joint sampling and routing policy $\pi\triangleq(R_{0},Z_{0},R_{1},Z_{1}\dots)$. This policy consists of two distinct sequences: (i) a sequence of routing decisions $(R_{0},R_{1},R_{2},\dots)$, where $R_{i}$ specifies the route selected for transmitting the $(i+1)$-th packet, and (ii) a sequence of finite waiting times $(Z_{0},Z_{1},Z_{2},\dots)$, where $Z_{i}<\infty$ determines the $(i+1)$-th sampling (or submission) time as $S_{i+1}=D_{i}+Z_{i}$.
Let $\Pi$ denote the set of all causal and stationary policies $\pi$. We consider the following problem:
Problem 1 (Average Age Penalty Minimization with Energy Constraint).
$\lambda^{\star}=$ $\min_{\pi\in\Pi}\;\limsup_{T\to\infty}\;\frac{1}{T}\mathbb{E}\!\left[\int_{0}^{T}f(\Delta(t))\,dt\right]$ (5)
@@
where $N_{s}(T)$ is the number of sampling actions up to time $T$, and $\lambda^{\star}$ is the optimal long-term average AoI cost. This problem aims at minimizing the long-term average age penalty under a long-term average energy constraint.
@@
$\limsup_{T\to\infty}\mathbb{E}\!\left[\frac{1}{T}\int_{0}^{T}\!f(\Delta(t))\,dt\right]=$ (6)
$\limsup_{n\to\infty}\frac{\mathbb{E}\!\left[\frac{1}{n}\sum_{i=0}^{n-1}q(Y_{i},Z_{i},Y_{i+1})\right]}{\mathbb{E}\!\left[\frac{1}{n}\sum_{i=0}^{n-1}(Z_{i}+Y_{i+1})\right]},$
where $q(Y_{i},Z_{i},Y_{i+1})$ represents the accumulated age penalty (the area under the curve $f(\Delta(t))$) during the $i$-th cycle. An illustrative example of this process for an exponential penalty function is shown in Fig. 3, where the shaded region corresponds to the accumulated cost $q(\cdot)$.
In addition, the energy process $E(t)$ is also a piecewise constant function. Similarly, the long-term average energy consumed up to the $n$-th update can be written as:
@@
$=\limsup_{n\to\infty}\frac{\mathbb{E}\!\left[\frac{1}{n}\sum_{i=0}^{n-1}E_{i+1}\right]}{\mathbb{E}\!\left[\frac{1}{n}\sum_{i=0}^{n-1}(Z_{i}+Y_{i+1})\right]},$
@@
Problem 2 (Reformulated Non-linear Age Minimization).
$\lambda^{\star}\triangleq\min_{\pi}\;$ $\limsup_{n\to\infty}\;\frac{\mathbb{E}\!\left[\frac{1}{n}\sum_{i=0}^{n-1}q(Y_{i},Z_{i},Y_{i+1})\right]}{\mathbb{E}\!\left[\frac{1}{n}\sum_{i=0}^{n-1}(Z_{i}+Y_{i+1})\right]}$ (8a)
s.t. $\limsup_{n\to\infty}\;\frac{\mathbb{E}\!\left[\frac{1}{n}\sum_{i=0}^{n-1}E_{i+1}\right]}{\mathbb{E}\!\left[\frac{1}{n}\sum_{i=0}^{n-1}(Z_{i}+Y_{i+1})\right]}\;\leq\;E_{\max}$ (8b)
Problem 2 is an infinite-horizon Constrained Semi-Markov Decision Process (CSMDP) with a hybrid state space and a state-dependent hybrid action space, posing significant challenges due to its fractional objective and the presence of a long-term energy constraint. To tackle this, we adopt a divide-then-concur strategy, consisting of three key steps: $i$) Fractional constraint reformulation: We first transform the original fractional constraint into an equivalent linear form to simplify the constraint structure; $ii$) Dinkelbach transformation: Next, we apply the Dinkelbach approach [12] to convert the fractional objective into a sequence of parameterized CMDP problems; iiiiii) Lagrangian techniques: Finally, we employ the Lagrangian method to handle the constraint in the Constrained Markov Decision Process (CMDP) formulation, effectively reducing the problem to a standard MDP.